OpenStudy (anonymous):

write the equation of the line, in slope intercept form, through (2,-4) and perpendicular to x-2y=7

OpenStudy (anonymous):

y = -2x

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

hmm perpendicular slope will be -2

OpenStudy (anonymous):

y = -2x + c determine c by putting in the point (2,-4)

myininaya (myininaya):

line has form y=mx+b we know y=mx+b is perpendicular to x-2y=7 what is the slope of x-2y=7? solve x-2y=7 for y x-7=2y y=x/2-7/2 slope is 1/2 so line perpendicular to y=x/2-7/2 would have slope -2 so now we know we have a line in the form y=-2x+b but we also want to know what line goes through (2,-4) so we plug that into find b -4=-2(2)+b -4=-4+b add 4 to both sides 0=b so the line is y=-2x

OpenStudy (anonymous):

Perfect !!